4. Find the angle BO1O2 when the links containing the sliding pair become vertical. 2 m 3 m 5 m 4 m 3 m
Answer #1

Draw the schematic diagram when the sliding pair link is vertical.

2 m B AO 모 4 m 5 m 900 02. 3 m

Consider triangle AO1O2,.

This is right angle triangle.

(011) = 52 - 32 ОА1 = 4 m

Now, consider triangle AO1B

As per cosine law of triangle.

(AB)2 = (101)2 + (OB) – 2 X A01 X 01B x cosa

substitute the respective values from the figure.

22 = 42 +42 – 2 x 4 x 4 x COSA cos @ 16 + 16 - 4 32 8 = cos-1 (0.875) A = 28.95°

From the figure,

0 + a = 90°

substitute A= 28.95° in the above equation.

28.95° + a = 90 a = 61.05°

Therefore, angle BO1O2 = 61.05%